# Model-oriented Graph Distances via Partially Ordered Sets

### v1

@@ G.3 Pseudo-distance upper-bounding SHD: proof of Proposition 13
Consider a pair of vertices $(v_{i},v_{j})$ such that the edge between them differs in $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. Then, based on the observation above, for any path $p\in L_{\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$, there exists at least one move along the path between graphs $\mathcal{G}_{x}$ and $\mathcal{G}_{y}$ where only this edge changes. Note that if the edge is missing in one graph and undirected in the other, or if the edge is directed in opposite orientations in the two graphs, then at least two such moves exist in the path. This implies that $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is lower bounded by the structural Hamming distance. ∎
G.4 Simplifications under polytree assumptions: proof of Theorem 4
The proof relies on the following lemma, which we prove first. Throughout, we note that an MPDAG $\mathcal{G}$ can be viewed as containing all DAGs that are members of a CPDAG $\mathcal{C}$ subject to a set of fixed directed edges.
Lemma 8.
Let $\mathcal{G}$ be a MPDAG that is a polytree and represents the collection of DAGs $\mathcal{D}_{1},\dots,\mathcal{D}_{k}$. Consider any pair of nodes $(i,j)$ that are connected in $\mathcal{G}$ and let $\tilde{\mathcal{D}}_{l}$ be the DAG obtained by removing the edge between nodes $(i,j)$ in $\mathcal{D}_{l}$. Then, the collection $\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2}\dots,\tilde{\mathcal{D}}_{k}$ forms a MPDAG $\tilde{\mathcal{G}}$ with $\tilde{\mathcal{G}}\preceq\mathcal{G}$.
Proof of Lemma˜8.
Since $\mathcal{G}$ is a valid MPDAG, all the DAGs it represents have the same skeleton and v-structures. By construction, all modified DAGs (after removing the edge) have the same skeleton. Further, they all have the same v-structure. In particular, each $\mathcal{D}_{i}$ is a polytree, and removing an edge does not uncover a v-structure. Further, if the edge removal removes a v-structure in one modified graph, it removes it for all the other modified graphs. We have thus shown that the collection of DAGs $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$ are all Markov equivalent and members of a CPDAG $\tilde{\mathcal{C}}$.
It remains to check that the collection $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$ forms a valid MPDAG. That is, the collection represents all DAGs that are members of $\tilde{\mathcal{C}}$, subject to a set of fixed directed edges $\tilde{\mathcal{E}}=\{i_{1}\to i_{2},i_{3}\to i_{4},\dots\}$. Let $\mathcal{C}$ be the CPDAG associated with the collection $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{k}\}$. By definition, $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{k}\}$ is all DAGs that are members of $\mathcal{C}$}, subject to a set of fixed directed edges ${\mathcal{E}}=\{j_{1}\to j_{2},j_{3}\to j_{4},\dots\}$. Let
$\tilde{\mathcal{E}}=\mathcal{E}\cup\{k\to j:(i,j,k)\in\operatorname{v-str}(\mathcal{C})\}\cup\{k\to i:(j,i,k)\in\operatorname{v-str}(\mathcal{C})\}.$
Let $\tilde{\mathcal{G}}$ be the MPDAG obtained by restricting $\tilde{\mathcal{C}}$ to the edge set $\tilde{\mathcal{E}}$. We will show that the collection of DAGs represented by $\tilde{\mathcal{\mathcal{G}}}$ is precisely $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$. First, we show the collection of DAGs represented by $\tilde{\mathcal{\mathcal{G}}}$ is a subset of $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$. Take a DAG $\tilde{\mathcal{D}}$ represented by $\tilde{\mathcal{G}}$. By construction, $\tilde{\mathcal{D}}$ is compatible with edge set $\mathcal{E}$ (since $\mathcal{E}\subseteq\tilde{\mathcal{E}}$). Add the edge between nodes $(i,j)$ in the same orientation as any DAG in $\mathcal{\mathcal{G}}$ to obtain a DAG $\bar{\mathcal{D}}$. Then, by construction, $\bar{\mathcal{D}}$ is a member of $\mathcal{C}$, and is compatible with the edge set $\mathcal{E}$. Thus, $\bar{\mathcal{D}}$ is among the collection $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{k}\}$, which means that $\tilde{\mathcal{D}}$ is among the collection $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$. Next, we show the collection of DAGs represented by $\tilde{\mathcal{\mathcal{G}}}$ is a superset of $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$. Recall that $\tilde{\mathcal{D}}_{i}$ is a member of $\tilde{\mathcal{C}}$ for every $i=1,2,\dots,k$. Furthermore, by construction, each $\tilde{\mathcal{D}}_{i}$ is compatible with $\tilde{\mathcal{E}}$. This direction then follows.
Proof of Theorem 4.
Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be two MPDAGs that are polytrees with $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$. Suppose as a point of contradiction that $\mathrm{rank}(\mathcal{G}_{1})+1<\mathrm{rank}(\mathcal{G}_{2})$ and there does not exist a distinct MPDAG that is a polytree in between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. We consider three cases and show that there is a contradiction in each case.
Case I: Two or more pairs of nodes are connected in $\mathcal{G}_{2}$ but not in $\mathcal{G}_{1}$. Let $(i,j)$ be one such pair. Let $\{\mathcal{D}_{1},\dots,\mathcal{D}_{k}\}$ be the collection of DAGs represented by $\mathcal{G}_{2}$. Let $\{\tilde{\mathcal{D}}_{1},\dots,\tilde{\mathcal{D}}_{k}\}$ to be the collection of DAGs obtained by removing the edge between pair $(i,j)$ from each $\mathcal{D}_{l},l=1,2,\dots,k$. By Lemma 8, the collection form an MPDAG $\tilde{\mathcal{G}}$ with $\tilde{\mathcal{\mathcal{G}}}\preceq\mathcal{G}_{2}$. Since $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$, we have that for every DAG $\mathcal{D}$ represented by $\mathcal{G}_{1}$, there exists a DAG ${\mathcal{D}}_{l}$ represented by $\mathcal{G}_{2}$ where $\mathcal{D}$ is a subgraph of ${\mathcal{D}}_{l}$. It then follows that $\mathcal{D}$ is also a subgraph of ${\mathcal{D}}_{l}$. So $\mathcal{G}_{1}\preceq\tilde{\mathcal{G}}$.
Case II: There is at least one pair of node that is connected via an undirected edge in $\mathcal{G}_{2}$ but is connected via a directed edge in $\mathcal{G}_{1}$, and there is one pair of nodes that are connected in $\mathcal{G}_{2}$ but not in $\mathcal{G}_{1}$. Suppose $\mathcal{G}_{2}$ is the MPDAG obtained by restricting the CPDAG $\mathcal{C}$ to edge set $\mathcal{E}$. Consider one of the undirected edges between nodes $(i,j)$ that is directed as $i\to j$ in $\mathcal{G}_{1}$. Let $\tilde{\mathcal{G}}$ be the MPDAG obtained by restricting $\mathcal{C}$ to the edge set $\mathcal{E}\cup\{i\to j\}$. Then, by construction, $\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$ with $\tilde{\mathcal{G}}\neq\mathcal{G}_{2}$. Furthermore, since $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$, for any DAG $\mathcal{D}_{1}$ represented by $\mathcal{G}_{1}$, there exists a DAG $\mathcal{D}_{2}$ represented by $\mathcal{G}_{2}$ such that $\mathcal{D}_{1}$ is a sub-graph of $\mathcal{D}_{2}$. Since $i\to j$ in $\mathcal{D}_{1}$, then $i\to j$ in $\mathcal{D}_{2}$. This implies that $\mathcal{D}_{2}$ is also represented by $\tilde{G}$, which consequently implies that $\mathcal{G}_{1}\preceq\tilde{\mathcal{G}}$. Since $\tilde{\mathcal{G}}$ has still at least one more edge than $\mathcal{G}_{1}$, we have that $\mathcal{G}_{1}\neq\tilde{\mathcal{G}}$.
Case III: There is at least two pair of node that are connected via an undirected edge in $\mathcal{G}_{2}$ but are connected via a directed edge in $\mathcal{G}_{1}$. Suppose $\mathcal{G}_{2}$ is the MPDAG obtained by restricting CPDAG $\mathcal{C}$ to edge set $\mathcal{E}$. Let $\tilde{\mathcal{G}}$ be the MPDAG obtained by restricting $\mathcal{C}$ to $\mathcal{E}\cup\{i\to j\}$. Following the argument as Case II, we have that $\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$ and $\mathcal{G}_{1}\preceq\tilde{\mathcal{G}}$. By construction, $\tilde{\mathcal{G}}\neq\mathcal{G}_{2}$. It remains to show that $\tilde{\mathcal{G}}\neq\mathcal{G}_{1}$.
Note that each pair of nodes that have a directed edge in $\mathcal{G}_{1}$ (but undirected in $\mathcal{G}_{2}$) cannot be part of a v-structure in $\mathcal{G}_{1}$. Otherwise, since $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are both polytrees, all the v-structures in $\mathcal{G}_{1}$ must also be present in $\mathcal{G}_{2}$, and so these edges would also need to be directed in $\mathcal{G}_{2}$.
We consider some sub-cases. In the first subcase, the two undirected edges are incident with one edge being between $(i,j)$ and the other between $(j,k)$. In this case, to avoid creating a v-structure, because of the polytree condition, $j\to k$ in $\mathcal{G}_{1}$. Suppose we let $\tilde{\mathcal{E}}=\mathcal{E}\cup\{j\to k\}$, and let $\tilde{\mathcal{G}}$ be the MPDAG obtained by restricting $\mathcal{C}$ to the edge set $\tilde{\mathcal{E}}$. Since the edge between $(i,j)$ is undirected in $\mathcal{G}_{2}$, we have that $\tilde{\mathcal{G}}$ has an undirected edge between node pair $(i,j)$. Thus $\tilde{\mathcal{G}}\neq\mathcal{G}_{1}$ in this subcase.
In the second subcase, the two undirected edges are incident with one edge being between $(i,j)$ and the other between $(i,k)$. Since the edge between $(i,k)$ is undirected in $\mathcal{G}_{2}$, we have that there must exist DAGs represented by $\mathcal{G}_{2}$ with the edge $k\to i$ and an edge $i\to j$ as well as DAGs with edge $i\to k$ and $i\to j$, as these do not create v-structures and there is no concern about violating acyclicity due to the polytree condition. We thus have that that $\tilde{\mathcal{G}}$ has an undirected edge between node pair $(i,k)$, and consequently $\tilde{\mathcal{G}}\neq\mathcal{G}_{1}$.
In the final subcase, the two undirected edges in $\mathcal{G}_{2}$ that are directed in $\mathcal{G}_{1}$ are not incident, i.e., they involve pairs $(i,j)$ and $(l,m)$. To be distinct from the previous subcases, together with the fact that $i\to j$ in $\mathcal{G}_{1}$ should not create a v-structure, means that all other edges involving $j$ must be of the form $j\to k$ in $\mathcal{G}_{1}$. This fact isolates the two edges between $(i,j)$ and $(l,m)$, meaning that taking among all DAGs represented by $\mathcal{G}_{2}$ that have the edge $i\to j$, there exists ones that have the edge $l\to m$ and ones that have the edge $m\to l$. Thus, the intermediate graph $\tilde{\mathcal{G}}$ is distinct from $\mathcal{G}_{1}$. ∎
Appendix H On the Idiosyncrasies of the structural interventional distance and separation distance
@@ H.1 Structural intervention distance
Figure 12: Three DAGs for which symmetrized SID violates the triangle inequality as shown in Example 2.
Example 2 (Symmetrized SID violates triangle inequality).
Let $\mathcal{G}_{1}$, $\mathcal{G}_{2}$, and $\mathcal{G}_{3}$ be the three DAGs from Figure 12. Denote $d_{SID}$, the structural intervention distance for DAGs as defined by (Peters and Bühlmann, 2015). Using the open source implementation for computing SID by (Henckel, Würtzen and Weichwald, 2024) we obtain that $d_{SID}(\mathcal{G}_{1},\mathcal{G}_{2})=3,d_{SID}(\mathcal{G}_{2},\mathcal{G}_{1})=4,d_{SID}(\mathcal{G}_{1},\mathcal{G}_{3})=5,d_{SID}(\mathcal{G}_{2},\mathcal{G}_{3})=5,d_{SID}(\mathcal{G}_{2},\mathcal{G}_{3})=9$ and $d_{SID}(\mathcal{G}_{3},\mathcal{G}_{2})=9$. For the symmetrized version of SID, denoted $d_{\mathrm{sySID}}$, it therefore holds that
@@ H.2 Separation distance
We interpret a probabilistic graphical model as the collection of all random variables satisfying the conditional independence constraints encoded by the graph. A closely related and popular alternative is to interpret a probabilistic graphical model directly as the collection of conditional independence statements encoded by the graph or more formally the set of all triples $(X,Y,Z)$ corresponding to a conditional independence statement of the form $X\perp\!\!\!\!\perp Y\mid Z$ (Studenỳ, 1992). Based on this interpretation of a CPDAG as a collection of triples (Wahl and Runge, 2025) have proposed an alternative distance notion called the separation distance. It defines the distance between two CPDAGs as the size of the symmetric difference of the two corresponding sets of triples. Similarly to our class of distances, the separation distance is a metric and model oriented, albeit on a different model interpretation. It can also be generalized to any class of probabilistic graphical models that encode only conditional independence statements. In fact any separation distance can be viewed as a minimal path distance on a Hasse diagram: Consider the set of all triples $X,Y,Z\subseteq\{1,\dots,n\}$ and the poset structure induced by subset inclusion on this set. This poset forms a lattice, with the meet and join between two sets of triples defined as their intersection and union, respectively. As a result, the minimal path distance on the corresponding Hasse diagram equals the down-up distance, which in turn equals the size of the symmetric difference between the two sets.
@@
Example 3.
Let $\mathcal{G}_{1}$ denote the empty CPDAG, $\mathcal{G}_{2}$ a CPDAG with a single (undirected) edge, $\mathcal{G}_{3}$ the fully connected CPDAG, and $\mathcal{G}_{4}$ the same graph with one (undirected) edge removed, all on $n$ nodes. Since all d-separation statements hold in $\mathcal{G}_{1}$, it can be represented by the set of all triples of subsets on the set of $n$ nodes. In $\mathcal{G}_{2}$, on the other hand, all statements hold except those of the form
$\mathrm{V}_{i}\perp_{\mathcal{G}_{2}}\mathrm{V}_{j}\mid Z,\quad Z\subseteq V\setminus\{\mathrm{V}_{i},\mathrm{V}_{j}\}.$
As a result, the separation distance between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ is the size of this latter set, that is, $2^{n-1}$. Furthermore, for every disconnected node we add to both $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, their distance doubles. The CPDAG $\mathcal{G}_{3}$, on the other hand, encodes no d-separation statements and therefore corresponds to the empty set, while $\mathcal{G}_{4}$ encodes the single statement $\mathrm{V}_{i}\perp\mathrm{V}_{j}\mid\mathrm{V}\setminus\{\mathrm{V}_{i},\mathrm{V}_{j}\}$. Hence, the separation distance between $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$ equals $1$. Again adding a disconnected node to both graphs doubles this distance.
@@
Appendix I Algorithmic Aspects of Computing the $\mathcal{M}$-oriented distance
We present additional algorithms that are important for finding the $\mathcal{M}$-oriented distance in the context of CPDAGs.
Up-down distance: The algorithm for computing the up-down distance is the ‘dual’ of Algorithm 2. In particular, (i) the neighbor selection step is restricted to be upward; (ii) the search terminates when a graph $\mathcal{G}^{\prime}$ is found with $\mathcal{G}_{t}\preceq\mathcal{G}^{\prime}$; (iii) we use the admissible heuristic $2|\mathrm{sk}(\mathcal{G}_{t})\cup\mathrm{sk}(\mathcal{G}^{\prime})|-\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}^{\prime})$; and finally (iv) we verify whether $\mathrm{rank}(\mathcal{G}^{\prime})\geq u:=|\mathrm{sk}(\mathcal{G}_{t})\cup\mathrm{sk}(\mathcal{G}^{\prime})|$ before checking if $\mathcal{G}_{t}\preceq\mathcal{G}^{\prime}$. Here, $u$ serves as the upper bound for $\rho_{\mathrm{join}}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min_{\mathcal{G}\in\mathfrak{G},\mathcal{G}_{s}\preceq\mathcal{G},\mathcal{G}_{t}\preceq\mathcal{G}}\mathrm{rank}(\mathcal{G})$; see Lemma 6 in Section F.3 for proof. Algorithm 3 presents the full description of the procedure for finding the up-down distance.
input : Graph class of CPDAGs $\mathfrak{G}$, pair of graphs $\mathcal{G}_{s},\mathcal{G}_{t}$, a heuristic $h\colon\mathfrak{G}\to\mathbb{R}_{\geq 0}$, upwards neighbor function $un_{\mathcal{L}}\colon\mathfrak{G}\to 2^{\mathfrak{G}}$, upper bound $u$ to $\rho_{\text{join}}(\mathcal{G}_{s},\mathcal{G}_{t})$
output : Up-down distance $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})$
1 begin
 $openSet\coloneqq\{\mathcal{G}_{s}\}$;
 // priority queue ordered by $g+h$
2 $g(\mathcal{G}_{s})\coloneqq 0$;
3 while $openSet\neq\emptyset$ do
4    $\mathcal{G}_{\text{current}}\coloneqq$ graph in openSetopenSet minimizing $g(\mathcal{G}_{\text{current}})+h(\mathcal{G}_{\text{current}})$;
5    if $u\leq\mathrm{rank}(\mathcal{G}_{\text{current}})\,\,\land\,\,\mathcal{G}_{t}\preceq\mathcal{G}_{\text{current}}$ then
6       return $g(\mathcal{G}_{\text{current}})+\mathrm{rank}(\mathcal{G}_{\text{current}})-\mathrm{rank}(\mathcal{G}_{t})$
7     remove $\mathcal{G}_{\text{current}}$ from openSetopenSet;
8
9    foreach $\mathcal{G}_{\text{neighbor}}\in$ un($\mathcal{G}_{\text{current}}$) do
10       $tentative\coloneqq g(\mathcal{G}_{\text{current}})+1$;

### v2

@@ G.2 Pseudo-distance upper-bounding SHD: proof of Proposition 11
Consider a pair of vertices such that the edge between them differs in $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. Then, based on the observation above, for any path $p\in\mathrm{P}_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$, then $p$ must go through some neighboring graphs $\mathcal{G}_{i},\mathcal{G}_{i+1}$ such that they differ exactly at this edge. If the edge is missing in one graph and undirected in the other, or if the edge is directed in opposite orientations in the two graphs, then at least two such moves exist in the path. This implies that $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is lower bounded by $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})$. ∎
Appendix H Idiosyncrasies of the structural interventional distance and separation distance
@@ H.1 Structural intervention distance
Figure S13: Three DAGs for which symmetrized SID violates the triangle inequality as shown in Example S3.
Example S3 (Symmetrized SID violates triangle inequality).
Let $\mathcal{G}_{1}$, $\mathcal{G}_{2}$, and $\mathcal{G}_{3}$ be the three DAGs from Figure S13. Denote $d_{SID}$, the structural intervention distance for DAGs as defined by (Peters and Bühlmann, 2015). Using the open source implementation for computing SID by (Henckel, Würtzen and Weichwald, 2024) we obtain that $d_{SID}(\mathcal{G}_{1},\mathcal{G}_{2})=3,d_{SID}(\mathcal{G}_{2},\mathcal{G}_{1})=4,d_{SID}(\mathcal{G}_{1},\mathcal{G}_{3})=5,d_{SID}(\mathcal{G}_{2},\mathcal{G}_{3})=5,d_{SID}(\mathcal{G}_{2},\mathcal{G}_{3})=9$ and $d_{SID}(\mathcal{G}_{3},\mathcal{G}_{2})=9$. For the symmetrized version of SID, denoted $d_{\mathrm{sySID}}$, it therefore holds that
@@ H.2 Separation distance
[…]
@@
Example S4.
Let $\mathcal{G}_{1}$ denote the empty CPDAG, $\mathcal{G}_{2}$ a CPDAG with a single (undirected) edge, $\mathcal{G}_{3}$ the fully connected CPDAG, and $\mathcal{G}_{4}$ the same graph with one (undirected) edge removed, all on $n$ vertices. Since all d-separation statements hold in $\mathcal{G}_{1}$, it can be represented by the set of all triples of subsets on the set of $n$ vertices. In $\mathcal{G}_{2}$, on the other hand, all statements hold except those of the form
[…]
@@
Appendix I Algorithms and additional numerical results
I.1 Proof of Theorem˜8
Proof.
Because $\mathcal{L}$ is connected and $\mathfrak{G}$ is finite, $d_{\mathcal{L}}$ is well-defined for every pair of graphs in $\mathfrak{G}$. Suppose the algorithm does not terminate. Then, it must be due to graphs continuing to be pushed to openSet in line 22. However, this means line 19 of the algorithm is executed infinitely many times, where after each time $g[\mathcal{G}^{\prime}]$ is either created or strictly decreased. We have a contradiction because this implies the condition in line 18 must fail at some point.
In what follows, we show that the algorithm returns the correct distance. Observe that the while loop cannot exit without returning either at line 8 or 12, because if so $\mathcal{G}_{t}$ is never visited and that contradicts the connectedness of $\mathcal{L}$. We first show that if the algorithm returns at line 8, $u^{\ast}$ upon return equals the desired distance. To see this, we argue that there must exist a shortest path that is routed through a graph $\mathcal{G}^{\prime}$ in openSet for which it holds that $f(\mathcal{G}^{\prime})\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$. Specifically, consider any shortest path $p^{\ast}$ from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$. Because $\mathcal{G}_{t}$ is not yet reached, we have $p^{\ast}\cap\textsf{openSet}\neq\emptyset$ and we choose $\mathcal{G}^{\prime}$ to be the graph in this intersection that is pushed to openSet most recently. Indeed, for this graph on the frontier, we have $g[\mathcal{G}^{\prime}]=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}^{\prime})$ and hence $f(\mathcal{G}^{\prime})=g[\mathcal{G}^{\prime}]+h(\mathcal{G}^{\prime})\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}^{\prime})+d_{\mathcal{L}}(\mathcal{G}^{\prime},\mathcal{G}_{t})=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ by the admissibility of $h$. Then, by the property of the priority queue, it follows that $\textsf{f}\leq f(\mathcal{G}^{\prime})\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq u^{\ast}$. Meanwhile, we also have $\textsf{f}\geq u^{\ast}$ by line 7 so we conclude $\textsf{f}=u^{\ast}=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
Now we argue that g returned at line 12 equals $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$. Suppose this does not hold. Observe that $\textsf{g}=g[\mathcal{G}_{t}]$ when $\mathcal{G}_{t}$ was pushed into openSet and by construction $g[\mathcal{G}_{t}]\geq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ holds, so we must have $\textsf{g}>d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$. Let $p^{\ast}$ be any shortest path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$. Similar to the argument in the preceding paragraph, let $\mathcal{G}^{\ast}$ be the graph on $p^{\ast}$ that was last pushed to openSet right before $\mathcal{G}_{t}$ was popped at line 6. By construction of the algorithm and the fact that $p^{\ast}$ is a shortest path, we have $g[\mathcal{G}^{\ast}]=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}^{\ast})$. Further, using admissibility of $h$, we know
$f(\mathcal{G}^{\ast})=g[\mathcal{G}^{\ast}]+h(\mathcal{G}^{\ast})\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}^{\ast})+d_{\mathcal{L}}(\mathcal{G}^{\ast},\mathcal{G}_{t})=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t}).$
However, because the priority queue pops the item with the smallest index, we also have $f(\mathcal{G}^{\ast})\geq f(\mathcal{G}_{t})=\textsf{g}+0>d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ and hence a contradiction. ∎
I.2 Subroutines for computing the distance between CPDAGs
Algorithm˜2 computes the lower bound on the model-oriented distance between two CPDAGs given by Proposition˜6. Algorithm˜3 computes $m_{\text{op}}(\cdot)$ for a set of v-structure triplets.
Input: CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ over a vertex set $\mathrm{V}$
Require : Subroutines $\textnormal{{FindMinOps}}(\cdot)$ and $\textnormal{{DisjointTriplets}}(\cdot)$
Output: The lower bound on $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ given by Proposition˜6
1 $\mathcal{G}_{s,\operatorname{sk}}\leftarrow\operatorname{sk}(\mathcal{G}_{s})$
2 $\mathcal{G}_{t,\operatorname{sk}}\leftarrow\operatorname{sk}(\mathcal{G}_{t})$
$d\leftarrow|\mathcal{G}_{s,\operatorname{sk}}\triangle\mathcal{G}_{s,\operatorname{sk}}|$
// skeleton difference
$\mathrm{S}^{\ast}\leftarrow\{\}$
// $\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})$
3 foreach $a\in\mathrm{V}$ do
4 foreach $b\in\mathrm{V}$ do
5    foreach $c\in\mathrm{V}$ do
6       if $a=b$ or $a=c$ or $b=c$ then
7          Continue
8       if $(a,b)\in\mathcal{G}_{s,\operatorname{sk}}\cap\mathcal{G}_{t,\operatorname{sk}}$ and $(b,c)\in\mathcal{G}_{s,\operatorname{sk}}\cap\mathcal{G}_{t,\operatorname{sk}}$ and $(a,c)\notin\mathcal{G}_{s,\operatorname{sk}}\cup\mathcal{G}_{t,\operatorname{sk}}$ then
9          if $(a\rightarrow b\leftarrow c$ in $\mathcal{G}_{s})$ and $(a\rightarrow b\leftarrow c$ not in $\mathcal{G}_{t})$ then
10             $\textnormal{{Push}}(\mathrm{S}^{\ast},(a,b,c))$
11          else if $(a\rightarrow b\leftarrow c$ in $\mathcal{G}_{t})$ and $(a\rightarrow b\leftarrow c$ not in $\mathcal{G}_{s})$ then
12             $\textnormal{{Push}}(\mathrm{S}^{\ast},(a,b,c))$
13
14
15
16